# ToolSandbox: A Stateful, Conversational, Interactive Evaluation Benchmark for LLM Tool Use Capabilities

### v1

@@ A.2.1 Tool Augmentations
A.3 Message Bus
@@ A.4 User Role
We introduce the implementation details of User Role in this section. As is shown in Figure 7, user simulator prompts consists of 3 components:
@@
User simulation error rate are consistent across representative agent models, as shown in Table 5. A decrease of error rate for Mistral was caused by the fact that, open source models would even struggle with prompting for user input correctly. Considering the wide performance gap between open source and proprietary models in Table 4, this does not affect our conclusions for Table 4.
GPT4-o Claude-3-Opus Gemini-1.5-Pro Mistral-7B-Instruct
Hallucination 6.90±1.45 6.40±0.97 7.15±0.71 3.10±0.46
IF 1.11±0.84 1.38±0.69 0.92±0.49 0.63±0.37
Total Error 8.02±1.36 7.78±0.52 8.07±1.03 3.73±0.64
Table 5: Percentage of user simulation failures in each failure category for each agent model. Mean and std collected from 4 repeated trials.
@@ A.5 Agent Role
Prompts used for all agents is shown in Figure 8. The prompt is meant to be consistent for all agents, and does not leak specific information about the testing environment.
@@
A.6 Evaluation
@@ B.3 Tool Design
Specifically for State Dependencies, we designed 4 types of state dependencies. 44% of all tools are either tools that depend on these states, manipulate them, or inspect them. These include:
  • Cellular service: Tools that require cellular service (e.g., send_message) depend on this state to be true.
  • Wifi: All RapidAPI-backed tools (e.g., search_stock) depend on this state to be true.
  • Location Service: All tools that utilize current location (e.g., get_current_location) depend on this state to be true.
  • Low battery mode: All tools that turn on cellular service, wifi, and location service status depend on this state to be false, creating nested state dependency.
B.4 ToolSandbox Category Statistics
@@ Appendix D Additional Evaluation Results
D.1 Model Feature Comparison
@@
In a conversational, interactive tool-use workflow, the agent needs to be able to accept multiple rounds of user input, decide when to generate a tool call or respond to user, and consume a tool response to determine the next step. However, as shown in 7, open source models including Gorilla and Command-R are not capable of consuming tool responses. Because of this, they can theoretically produce reasonable results for Single Tool Call test scenarios, but cannot complete any test scenario that requires multiple tool calls.
D.2 Turn Count Comparison

### v2

@@ A.2 Tools
Tools in ToolSandbox are a set of highly composable, explicitly or implicitly dependent Python functions, creating complex reasoning challenges. Besides python native tools, a handful of carefully selected RapidAPI tools were also included with a thin layer of python wrapper. Tools manipulate world state through the Execution Context when necessary, and raise informative exceptions when execution conditions were not met. As an example, in Figure 1, when send_message tool is called while cellular service is off, a ConnectionError is raised. This allows the Agent to reason over possible exceptions, and deduce the tool needed to resolve the exception.
@@ A.2.1 Tool Augmentations
A.3 Roles and Message Bus
In ToolSandbox there are three roles: User, Agent (Assistant) and Execution Environment. The Execution Environment, as a dedicated role, is responsible for executing tool-use requests from the Agent and returning the results. Interaction between the roles is enabled through a message passing system. Each message contains a sender role, recipient role, content as well as to which roles the message is visible to. A simple orchestrator determines message passing order by allowing the most recent recipient to be the next sender. Instead of representing the conversation as a single message thread, we use a collection of messages, i.e. Message Bus, stored within the Execution Context. The Message Bus contains a linear history of message transactions between all roles. As is shown in Figure 4, each role writes to the same Message Bus. However, when reading from the Message Bus, each role can only access a sub-view of the Message Bus based on which roles are allowed to "see" the individual messages. We will introduce each role in the following paragraphs.
@@ A.4 User Role
As is shown in Figure 7, user simulator prompts consists of 3 components:
@@ A.5 Agent Role
Initially, the Agent role will receive a message from the User in natural language. The Agent could decide to prompt the User again for additional information, or decide to issue a tool call towards the Execution Environment. When issuing a tool call, the Agent selects the name of the tool from a list of available tools and provides necessary arguments, commonly expressed as JSON objects. These JSON objects are converted to executable Python code, see Figure 9, and sent to the Execution Environment for execution. Prompts used for all agents is shown in Figure 8. The prompt is meant to be consistent for all agents, and does not leak specific information about the testing environment.
@@
A.6 Execution Environment Role
Some LLMs support parallel tool calling, intended to increase efficiency when multiple, independent tools need to be called. For example, if the user asks for the weather in two cities, the queries to the weather API can happen in parallel. However, if an LLM uses parallel tool calls for dependent tools, it should be penalized accordingly. For example, in Figure 1 where the agent has to enable cellular service before sending a text message, parallel tool calls should not be used. Execution Environment handles race conditions in parallel tool calls by following Murphy’s Law, ensuring race condition always happens if detected.
A.7 Evaluation
@@ B.3 Tool Design
B.4 ToolSandbox Statistics
@@
The number of milestones defined per world state database and column name can be found in Table 7.
Database Column Associated Milestone Count
CONTACT is_self 48
CONTACT name 56
CONTACT person_id 136
CONTACT phone_number 88
CONTACT relationship 56
MESSAGING content 40
MESSAGING recipient_phone_number 40
REMINDER content 136
REMINDER latitude 40
REMINDER longitude 40
REMINDER reminder_id 32
REMINDER reminder_timestamp 152
SETTING cellular 56
SETTING location_service 56
SETTING low_battery_mode 120
SETTING wifi 136
Table 7: Number of milestones defined per world state database and column name. Note that one test scenario can define multiple milestones, with multiple database constraints each.
The number of milestones defined per tool can be found in Table 8.
Tool Associated Milestone Count
calculate_lat_lon_distance 32
convert_currency 16
datetime_info_to_timestamp 32
get_cellular_service_status 8
get_current_location 48
get_current_timestamp 296
get_wifi_status 8
search_contacts 168
search_holiday 56
search_lat_lon 24
search_location_around_lat_lon 48
search_messages 104
search_reminder 80
search_stock 24
search_weather_around_lat_lon 80
timestamp_diff 96
unit_conversion 56
Table 8: Number of milestones defined per tool. Note that one test scenario can define multiple milestones, with multiple tool trace constraints each. Most state modification tool calls are tracked by corresponding database milestones instead of tool trace.
@@ Appendix D Additional Evaluation Results
D.1 Prompting Experiment
We conducted additional experiments comparing the effect of ReAct prompting in tool use agents against the minimal prompting design presented in Figure 8, with 4 representative models shown below.
From Table 9, we can see that, ReAct has minimal effect to Claude model family, and a small gain for GPT4 model family. This is likely due to the fact that, the baseline ToolSandbox setup, which allows the model to interact with execution environment, recover from error and confirm with simulated user already provides a reasoning process to the model through user and execution environment interactions, additional natural language reasoning would only provide marginal gains. While ReAct provide gains in some cases, relative ranking between models still stands. This backs up our claim in Section 4, where we considered prompt engineering gains orthogonal to the innate model capability surfaced by simpler prompting.
Baseline Score ReAct Score
GPT-4o-2024-05-13 73.0 73.6
Claude-3-Opus-20240229 69.2 69.3
GPT-4-0125-Preview 64.3 65.2
Claude-3-Sonnet-20240229 63.8 63.7
Table 9: A comparison between default prompting and ReAct prompting.
D.2 Model Feature Comparison
@@
In a conversational, interactive tool-use workflow, the agent needs to be able to accept multiple rounds of user input, decide when to generate a tool call or respond to user, and consume a tool response to determine the next step. However, as shown in 10, open source models including Gorilla and Command-R are not capable of consuming tool responses. Because of this, they can theoretically produce reasonable results for Single Tool Call test scenarios, but cannot complete any test scenario that requires multiple tool calls.
D.3 Turn Count Comparison